# The Reasoning Lingua Franca: A Double-Edged Sword for Multilingual AI

### v1

@@ 1 Introduction
In this study, we evaluate the question on two benchmark datasets, MGSM (Shi et al., 2022) and GPQA Diamond, which vary in difficulty (Rein et al., 2024) and together provide a comprehensive assessment of LRMs’ reasoning capabilities. We analyze results from three perspectives: (1) which reasoning language yields correct final answers more frequently; (2) which reasoning sequences exhibit richer cognitive behaviors; and (3) in which conditions reasoning in the question’s language outperforms reasoning in English.
@@ 6 Analyzing cognitive behaviors in the reasoning trace
To capture this, we analyze reasoning traces enclosed in <think></think> tokens for cognitive attributes such as Sub-goal setting, Verification, Back-tracking, and Backward chaining (c.f. Appendix D). We use gpt-4o-mini as a judge to identify and count these attributes, employing evaluation prompts (c.f. Appendix D.1) adapted from prior studies (Gandhi et al., 2025). For the MGSM task, we compute the average frequency of each attribute for the Qwen QwQ-32B model across languages, and report means and standard deviations over four trials.
@@ 7 Making a case for reasoning in the Question’s language
To quantify “Lost in Translation,” we measure the fraction of incorrect answers from English reasoning caused by translation mistakes; errors that would not arise if reasoning occurred directly in the question’s language. These cases are identified using gpt-4o-mini as evaluator. The evaluation setup and prompts are described in Appendix E.1. We average the proportion across four trials with the Qwen QwQ-32B model and report the standard deviation. This averaged proportion is referred to as the LiT (Lost in Translation) fraction. We manually verified the evaluator LLM’s judgments on a subset of examples from this Section and Section 6 to confirm the correctness of its decisions.
@@ Appendix D Analyzing Cognitive Behaviors: Additional details
We examine the following cognitive behaviors in the reasoning thread of LRMs:
  • Sub-goal setting: breaking a task into smaller, manageable steps.
  • Verification: performing systematic checks to catch and correct errors.
  • Backtracking: abandoning unsuccessful strategies and selecting alternative approaches.
  • Backward chaining: reasoning from the desired outcome back toward the necessary initial inputs.
These behaviors reflect how expert practitioners tackle difficult problems; for example, a software engineer tests and reviews each change, rewrites code when a design fails, and splits large features into smaller, testable modules.
D.1 Analyzing cognitive behaviors - prompt template and setup
@@ E.1 Lost in Translation - Quantitative Analysis
We use gpt-4o-mini as an evaluator LLM, employing the prompt template shown in Table 5 to determine whether an incorrect answer from reasoning in English ocurred due to a mistranslation or not. From gpt-4o-mini’s output we extract a "Yes" or "No" enclosed in <classification></classification> tags, compute the fraction of incorrect answers attributable to mistranslation for each run, and report the mean value over four trials.

### v2

@@ 1 Introduction
In this study, we evaluate the above question on two benchmark datasets, MGSM (Shi et al., 2022) and GPQA Diamond, which vary in difficulty (Rein et al., 2024) and together provide a comprehensive assessment of LRMs’ reasoning capabilities. We analyze results from three perspectives: (1) which reasoning language yields correct final answers more frequently; (2) which reasoning sequences exhibit richer cognitive behaviors; and (3) in which conditions reasoning in the question’s language outperforms reasoning in English.
@@ 6 Analyzing cognitive behaviors in the reasoning trace
To capture this, following prior studies (Gandhi et al., 2025), we analyze reasoning traces enclosed in <think></think> tokens for cognitive attributes such as Sub-goal setting (breaking a task into smaller, manageable steps), Verification (performing systematic checks to catch and correct errors), Back-tracking (abandoning unsuccessful strategies and selecting alternative approaches) , and Backward chaining (reasoning from the desired outcome back toward the necessary initial inputs). These behaviors reflect how expert practitioners tackle difficult problems; for example, a software engineer tests and reviews each change, rewrites code when a design fails, and splits large features into smaller, testable modules.
We use gpt-4o-mini as a judge to identify and count these attributes, employing evaluation prompts (c.f. Appendix D) adapted from prior studies (Gandhi et al., 2025). For the MGSM task, we compute the average frequency of each attribute for the Qwen QwQ-32B model across languages, and report means and standard deviations over four trials.
@@ 7 Making a case for reasoning in the Question’s language
To quantify “Lost in Translation,” we measure the fraction of incorrect answers from English reasoning caused by translation mistakes; errors that would not arise if reasoning occurred directly in the question’s language. These cases are identified using gpt-4o-mini as evaluator. The evaluation setup and prompts are described in Appendix E.1. We average the proportion across four trials with the Qwen QwQ-32B model and report the standard deviation. This averaged proportion is referred to as the LiT (Lost in Translation) fraction. Appendix E.2 details the human-LM agreement study for the Lost in Translation (LiT) task.
@@ 10 Ethics Statement
11 Acknowledgments
We thank AI4Bharat lab members Deepon Halder, Irina Agastin, and Atharva Gawane for their assistance with annotation on the Lost in Translation task. We would also like to thank EkStep Foundation and Nilekani Philanthropies for their generous grant towards research at AI4Bharat.
@@ E.1 Lost in Translation - Quantitative Analysis
We use gpt-4o-mini as an evaluator LLM, employing the prompt template shown in Table 6 to determine whether an incorrect answer from reasoning in English occurred due to a mistranslation or not. From gpt-4o-mini’s output we extract a "Yes" or "No" enclosed in <classification></classification> tags, compute the fraction of incorrect answers attributable to mistranslation for each run, and report the mean value over four trials.
E.2 Lost in Translation: Human-LM agreement
To validate the LLM-as-a-judge evaluator, we conducted a human annotation study for "Lost in Translation" (LiT) task. We randomly sampled 50 LiT evaluation prompts from MGSM for each of three low-resource languages - Hindi, Bengali and Malayalam, yielding 150 examples in total. Each example was independently solved and labeled by native-speaker annotators with strong grade-school mathematics skills. All annotators were computer science undergraduate students and each annotated 50 questions in their respective native language. Annotators were instructed to classify a reasoning trace as “Lost in Translation” or not based on the same instructions provided to the LLM judge. We compared the LLM judgments to the human labels and computed percent agreement, precision, recall, and Cohen’s K which are detailed in Table 5. These results indicate that the LLM evaluator produces reasonably reliable annotations for the LiT task even for low resource languages like Malayalam.
Language (Samples) Percent agreement Precision Recall Cohen’s kappa
Hindi (50) 0.78 0.83 0.80 0.55
Malayalam (50) 0.76 0.79 0.79 0.50
Bengali (50) 0.78 0.72 0.82 0.56
Total (150) 0.77 0.78 0.80 0.54
Table 5: Lost in Translation: Human–LLM agreement across 150 samples (50 per language).